# Intrinsic Second-Order magnon Thermal Hall Effect

### v1

@@ Abstract
In this paper, we study the intrinsic contribution of nonlinear magnon thermal Hall Effect. We derive the intrinsic second-order thermal Hall conductivity of magnon by the thermal scalar potential (TSP) method and the thermal vector potential (TVP) method. We find that the intrinsic second order magnon thermal Hall conductivity is related to the thermal Berry-connection polarizability (TBCP). We apply our theory to the monolayer ferromagnetic Hexagonal lattice, and we find that the second order magnon thermal Hall conductivity can be controlled by changing Dzyaloshinskii-Moriya strength and applying strain.
@@ I Introduction
The theory of electron transport is very important for the development of new electronic devices. With the improvement of wave packet dynamics, quantum transport theory of electrons has been developed a series of developments ref1 ; ref2 ; ref3 ; ref4 ; ref5 ; ref6 .
@@
For ferromagnets and antiferromagnets, the ground state of the system is usually an ordered state with ordered spin arrangement of magnetic ions. When some perturbation deviates the spin on a lattice from the direction of the quantized axis, the spins on neighboring lattices will be deviated as well, therefore the perturbation propagates in the form of collective motion, known as spin wave, with the quasiparticle named magnon. In particular, magnon can be excited and propagated in electronic insulators so that, in contrast to electrons, there is no energy dissipation due to the absence of joule heat. Therefore, magnons are promising to replace electrons as new information carriers. To this end, it is helpful to study the transport of magnons to develop functional devices based on new strategy. In order to regulate magnon transport in materials, various proposals have been studied extensively.
@@
In 2010, Katsura et al. predicted the magnon thermal Hall effect theoretically ref7 . In the same year, Onose et al. observed the magnon thermal Hall effect in ferromagnetic material with pyrochlorite lattice structure ref8 . After the work of Onose et al., the magnon thermal Hall effect and the magnon Hall-like effect have been found in different magnetic materials ref9 ; ref10 ; ref11 . Similar to the electron, magnon thermal transport is related to Berry curvature of magnon energy bands. In 2011, Matsumoto and Murakami studied the linear thermal Hall effect ref12 , and they found the relationship between linear magnon transport and Berry curvature. In 2013, Hoogdalem et al. demonstrated that the equilibrium magnetic texture leads to a fictitious magnetic field which can induce the magnon thermal Hall effect ref13 . The discovery of the magnon Hall effect has led to the study of the topological properties of magnon ref14 . In 2022, Kondo and Akagi proposed the nonlinear extrinsic magnon transport by Boltzmann equation ref15 . However, at present, the theory of intrinsic nonlinear transport of magnons is still lacking.
@@
In this work, we propose an intrinsic nonlinear magnon thermal Hall effect. Based on the work of Matsumoto and Murakami ref12 , by introducing the Berry curvature correction induced by temperature gradient, we obtain the intrinsic second-order transport of magnon. Then we take a numerical calculation of the two-dimensional ferromagnetic model to study the effect of strain and Dzyaloshinskii-Moriya interaction (DMI) on the intrinsic nonlinear magnon thermal Hall conductivity. We find that for Hexagonal ferromagnet (FM) to obtain non-zero intrinsic second-order Hall thermal conductivity of magnon, it is necessary to apply strain to break the symmetry of $\mathcal{C}_{3}^{z}$. In addition, more effects of strain on the intrinsic magnon second-order Hall thermal conductivity are also analyzed in this work.
@@ II Formalism
To describe the effect of temperature gradients in a Hamiltonian formalism, Luttinger introduced a fictitious scalar field $\Psi$, which is called the “gravitational” potential ref16 . It couples to energy density $h(\mathbf{r})$. To describe the effect of the temperature gradient, the Luttinger Hamiltonian reads
@@
Tatara introduced a thermal vector potential (TVP) in 2015 to describe the temperature gradient ref17 . In this way, the Luttinger Hamiltonian can alternatively be written as
@@
In analogy with the electron system, the dynamics of magnon wave packet can be described as ref12
@@ II.1.1 The perturbation from temperature gradients
For our purpose of time-dependent calculation, we follow the method in Ref. ref18, and introduce
@@
The Hamiltonian taking into account the effect of temperature gradient can be given by ref17 ; ref19
@@ II.2 The thermal Berry-connection polarizability and the correction of Berry curvature
The second-order nonlinear intrinsic Hall effect of electrons is related to Berry-connection polarizability (BCP), and the BCP is defined as $\mathcal{A}_{n,i}^{\prime}=G_{n,ji}E_{i}$, in which $E_{i}$ means the component of the electric field in the $i$ direction ref5 ; ref21 . The BCP can be expressed as
@@
in which subscripts $\alpha,\beta,\gamma$ and $\delta$ are Cartesian coordinates. Eqs. (21) and (25) are the one of central results of the present work.
@@ II.3 Intrinsic second-order magnon thermal Hall effect
Matsumoto and Murakami derived the linear magnon thermal Hall effect in two-dimensional system ref12 . The result is
@@
From Eq. (28) and keep to the second-order to the temperature gradient, we get the intrinsic second-order thermal Hall current of magnon as (see Appendix D)
@@
in which $i,j,\delta=x,y$ (see Appendix D). In the work of Liu et. al., the conductivity of the intrinsic nonlinear Hall effect is expressed as ref21
@@ III.1 Model
$+$ $g_{J}\mu_{B}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\mathbf{B}+K\sum_{i}\hat{S}_{i}^{z2}.$
The first term is Heisenberg interaction, the second term is the DM interaction, the third term is the Zeeman interaction, and the fourth term is the easy-axis anisotropy. For the symmetry of $\mathcal{C}_{3}^{z}$ and $\mathcal{P}$, 2THCV satisfies
@@
$\hat{\Phi}(\mathbf{k})=(\hat{a}_{\mathbf{k}},\hat{b}_{\mathbf{k}})^{T},\$
@@
Here, $\Delta_{BSa}=-3JS-g_{J}\mu_{B}B-2KS$, $\Delta_{BSb}=-3JS+g_{J}\mu_{B}B-2KS$, $J=\left(J_{1}+J_{2}+J_{3}\right)/3$, $\gamma_{\mathbf{k}}=\frac{1}{3J}\sum_{i}J_{i}e^{i\mathbf{k}\cdot\mathbf{\delta_{i}}}$, and $\Delta(\mathbf{k})=2SD\left[\sin{\frac{1}{2}(k_{y}-k_{x}\sqrt{3})}-\sin{k_{y}}+\sin{\frac{1}{2}(k_{y}+k_{x}\sqrt{3})}\right]$ (see Appendix E).
@@ III.2 Numerical result
To describe the effect of uniaxial strain along the direction of $\bm{\delta}_{2}$, we follow the method in Ref. ref15, in which $J_{1}=J_{3}=1$ meV is kept and $J_{2}$ running as a parameter. We use $J_{2}=J_{2}^{0}\delta$ where $J_{2}^{0}=1$ meV and $\delta$ takes into account the effect of strain on changing of $J_{2}$. And in our work, we take $g_{J}\mu_{B}B=-0.1$ meV, and $K=-0.2$ meV.
@@
Then we study the effects of strain and changing DM strength on $\kappa_{xyy}$. In Fig. 5 (a) and (b), we plot $\kappa_{xyy}$ curve with respect to temperature under different strain and different DM strength respectively. As the temperature rises, $\kappa_{xyy}$ increases rapidly at low temperature, and decreases after reaching a maximum value. There is no $\kappa_{xyy}$ without strain. $\kappa_{xyy}$ with strain of $\delta=0.8$ and $\delta=1.2$ have different sign. And the $\kappa_{xyy}$ curve increases with the increasing of DM strength. As shown in Fig. 5 (c), we can reverse the direction of intrinsic second-order magnon thermal Hall current by changing the direction of the strain. And we can increase the intrinsic second-order magnon Hall thermal conductivity by increasing the strength of strain. As shown in Fig. 5 (d), the intrinsic second-order magnon Hall thermal conductivity increases with the increasing of DM strength.
@@
When the second version of the present manuscript is being prepared, we note a similar work from Harsh Varshney et al. in which they calculated the density matrix modification of temperature gradient, then calculated the thermodynamic average value of magnon energy multiplied by velocity and the energy magnetization current through the density matrix ref22 . The intrinsic part of heat current has been extracted under the dilute impurity limit (DIL). In our work, we perform extra symmetry analysis in two-dimensional hexagonal ferromagnetic systems. We find that for existence of the non-zero intrinsic nonlinear Hall thermal conductivity of magnon in hexagonal ferromagnet, it is necessary to apply a strain to break the symmetry of $\mathcal{C}_{3}^{z}$. However, in the model calculation by Harsh Varshney et al., intrinsic nonlinear magnon thermal conductivity can be obtained in non-strain hexagonal ferromagnet model, which is different to our results. In addition, we propose the notion of thermal Berry-connection polarizability and work out compact formalism with transparent physical picture, which is out of the scale of Ref. ref22, . Moreover, we study the effects of strain on the second order intrinsic nonlinear thermal Hall transport.
@@ Appendix A The detail of the TVP perturbation method
We denote the zero-order perturbed part of $a_{m}(\mathbf{k},t)$ as $a_{m}^{(0)}(\mathbf{k},t)$, and denote the first-order perturbed part of $a_{m}(\mathbf{k},t)$ as $a_{m}^{(1)}(\mathbf{k},t)$. The initial condition is $a_{m}(\mathbf{k},0)=\delta_{mn}$, and the zeroth-order coefficients satisfy $\dot{a}_{m}^{(0)}(\mathbf{k},t)=0$. So we can get $a_{m}^{(0)}(\mathbf{k},t)=\delta_{mn}$. Substituting the zeroth-order coefficients into formula (S9), we can know that the first-order coefficients satisfy
@@
Then we take the zero frequency limit $\omega\to 0$, and remove non-physical items that increase over time ref18
@@
To get a steady-state, we should introduce a relaxation $\tau$ due to scattering. We take the long-time approximation $(t\gg\tau)$, such that all oscillatory terms drop out ref18 . Therefore, we can get the corrected Bloch state
@@ B.0.1 The perturbation from temperature gradient
Analogy the gravitational potential in electron system Smrcka1977 , Hamiltonian under the temperature gradient can be described as
@@ B.0.2 The corrected Bloch state
We follow the method in Ref. ref5, , and derive the modified Bloch state. Without the perturbation, the wave packet can be expressed as
@@
From ref2 , we know $\bra{\psi_{l}(\mathbf{k}^{\prime})}\hat{\mathbf{r}}\ket{\psi_{m}(\mathbf{k})}=i\partial_{\mathbf{k}^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})\delta_{lm}+\bm{\mathcal{A}}_{lm}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, so we can get

### v2

@@ Abstract
In this paper, we study the intrinsic contribution of nonlinear magnon thermal Hall Effect. We derive the intrinsic second-order thermal Hall conductivity of magnon by the thermal scalar potential (TSP) method and the thermal vector potential (TVP) method. We find that the intrinsic second-order magnon thermal Hall conductivity is related to the thermal Berry-connection polarizability (TBCP). We apply our theory to the monolayer ferromagnetic Hexagonal lattice, and we find that the second-order magnon thermal Hall conductivity can be controlled by changing Dzyaloshinskii-Moriya strength and applying strain.
@@ I Introduction
The theory of electron transport is very important for the development of new electronic devices. With the improvement of wave packet dynamics, quantum transport theory of electrons has been developed a series of developments [1, 2, 3, 4, 5, 6].
@@
For ferromagnets and antiferromagnets, the ground state of the system is usually an ordered state with ordered spin arrangement of magnetic ions. When some perturbation deviates the spin on a lattice from the direction of the quantized axis, the spins on neighboring lattices will be deviated as well, therefore the perturbation propagates in the form of collective motion, known as spin wave, with the quasiparticle named magnon. In particular, magnon can be excited and propagated in electronic insulators so that, in contrast to electrons, there is no energy dissipation due to the absence of joule heat [7]. Therefore, magnons are promising to replace electrons as new information carriers. To this end, it is helpful to study the transport of magnons to develop functional devices based on new strategy. In order to regulate magnon transport in materials, various proposals have been studied extensively.
@@
In 2010, Katsura et al. predicted the magnon thermal Hall effect theoretically [8]. In the same year, Onose et al. observed the magnon thermal Hall effect in ferromagnetic material with pyrochlorite lattice structure [9]. After the work of Onose et al., the magnon thermal Hall effect and the magnon Hall-like effect have been found in different magnetic materials [10, 11, 12, 13]. Similar to the electron, magnon thermal transport is related to Berry curvature of magnon energy bands. In 2011, Matsumoto and Murakami studied the linear thermal Hall effect [14], and they found the relationship between linear magnon transport and Berry curvature. The work of Murakami et al. initiated the study of the topology of magnon transport [15, 16, 17, 18]. In 2012, Onose et al. measured the Hall thermal conductivity of magnon in pyrochlore ferromagnetic insulators $Ho_{2}V_{2}O_{7}$ and $In_{2}Mn_{2}O_{7}$ [19]. And they found that the experimental results are in agreement with the theory proposed by Murakami et al. at low temperature. [14]. In 2013, Hoogdalem et al. demonstrated that the equilibrium magnetic texture leads to a fictitious magnetic field which can induce the magnon thermal Hall effect [20]. The discovery of the magnon Hall effect has led to the study of the topological properties of magnon [21]. In 2022, Kondo and Akagi proposed the extrinsic nonlinear magnon spin Nernst effect by Boltzmann equation [22]. In 2023, R. Mukherjee et al. proposed the extrinsic nonlinear thermal Hall effect of magnon by Boltzmann equation [23]. However, at present, the theory of intrinsic nonlinear transport of magnons is still lacking.
@@
In this work, we propose an intrinsic nonlinear magnon thermal Hall effect. Based on the work of Matsumoto and Murakami [14], by introducing the Berry curvature correction induced by temperature gradient, we obtain the intrinsic second-order transport of magnon. Then we take a numerical calculation of the two-dimensional ferromagnetic model to study the effect of strain and Dzyaloshinskii-Moriya interaction (DMI) on the intrinsic nonlinear magnon thermal Hall conductivity. We find that for hexagonal ferromagnet (FM) to obtain non-zero intrinsic second-order Hall thermal conductivity of magnon, it is necessary to apply strain to break the symmetry of $\mathcal{C}_{3}^{z}$. In addition, more effects of strain on the intrinsic magnon second-order Hall thermal conductivity are also analyzed in this work. When the first-order and second-order thermal Hall effect coexist, a non-reciprocal effect may be induced when reversing the temperature gradient. This is a kind of diode-like behavior which may be useful in future.
@@ II Formalism
To describe the effect of temperature gradients in a Hamiltonian formalism, Luttinger introduced a fictitious scalar field $\Psi$, which is called the “gravitational” potential [24]. It couples to energy density $h(\mathbf{r})$. To describe the effect of the temperature gradient, the Luttinger Hamiltonian reads
@@
The method of TSP often leads to unphysical divergent results as $T\to 0$. The divergent can be eliminated by the thermal vector potential (TVP) method introduced by Tatara [25, 26]. In this way, the Luttinger Hamiltonian can alternatively be written as
@@
In analogy with the electron system, the dynamics of magnon wave packet can be described as [14]
@@ II.1.1 The perturbation from temperature gradients
For our purpose of time-dependent calculation, we follow the method in the Appendix B of Ref. 27 and introduce
@@
The Hamiltonian taking into account the effect of temperature gradient can be given by [25, 26]
@@ II.2 The thermal Berry-connection polarizability and the correction of Berry curvature
The second-order nonlinear intrinsic Hall effect of electrons is related to Berry-connection polarizability (BCP), and the BCP is defined as $\mathcal{A}_{n,i}^{\prime}=G_{n,ji}E_{i}$, in which $E_{i}$ means the component of the electric field in the $i$ direction [5, 28]. The BCP can be expressed as
@@
in which subscripts $\alpha,\beta,\gamma$ and $\delta$ are Cartesian coordinates. Eqs. (21) and (25) are the one of central results of the present work. In short, when the temperature gradient is present, the thermal field (standing for the temperature gradient) will disturb the Berry connection, so as to the Berry curvature of the system. However, the thermal field $\mathbf{E}_{T}$ enters into the second-order heat current density rather than the second-order thermal Hall conductivity (see Eq. (29)).
@@ II.3 Intrinsic second-order magnon thermal Hall effect
Matsumoto and Murakami derived the linear magnon thermal Hall effect in two-dimensional system [14]. The result is
@@
From Eq. (28) and keep to the second-order to the temperature gradient, we get the intrinsic second-order Hall heat current of magnon as (see Appendix D)
@@
in which $i,j,\delta=x,y$ (see Appendix D). In the work of Liu et. al., the conductivity of the intrinsic nonlinear Hall effect is expressed as [28]
@@ III.1 Model
$+$ $g_{J}\mu_{B}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\mathbf{B}+\mathcal{K}\sum_{i}\hat{S}_{i}^{z2}.$
The first term is Heisenberg interaction, the second term is the DM interaction, the third term is the Zeeman interaction, and the fourth term is the easy-axis anisotropy. Eq. (36) formally contains the same terms as the Hamiltonian in the study about rectification of the spin Seebeck current by L. Chotorlishvili et al. [29]. However, Eq. (36) describes a two-dimensional hexagonal ferromagnetic lattice, and the Hamiltonian in the work of L. Chotorlishvili et al. describes the antiferromagnetic chain and two ferromagnetic chains with antiferromagnetic coupling [29].
For the symmetry of $\mathcal{C}_{3}^{z}$ and $\mathcal{P}$, 2THCV satisfies
@@
$\hat{\Phi}(\mathbf{k})=(\hat{a}_{\mathbf{k}},\hat{b}_{\mathbf{k}})^{T},\$ (39)
@@
Here, $\Delta_{BSa}=-3JS-g_{J}\mu_{B}B-2\mathcal{K}S$, $\Delta_{BSb}=-3JS+g_{J}\mu_{B}B-2\mathcal{K}S$, $J=\left(J_{1}+J_{2}+J_{3}\right)/3$, $\gamma_{\mathbf{k}}=\frac{1}{3J}\sum_{i}J_{i}e^{i\mathbf{k}\cdot\mathbf{\delta_{i}}}$, and $\Delta(\mathbf{k})=2SD\left[\sin{\frac{1}{2}(k_{y}-k_{x}\sqrt{3})}-\sin{k_{y}}+\sin{\frac{1}{2}(k_{y}+k_{x}\sqrt{3})}\right]$ (see Appendix E).
@@ III.2 Numerical result
To describe the effect of uniaxial strain along the direction of $\bm{\delta}_{2}$, we follow the method in Ref. 22 and Ref. 30 in which $J_{1}=J_{3}=-1$ meV is kept and $J_{2}$ running as a parameter. We use $J_{2}=J_{2}^{0}\delta$ where $J_{2}^{0}=-1$ meV and $\delta$ takes into account the effect of strain on changing of $J_{2}$. Here, $J_{2}$ is the Heisenberg spin exchange interaction with the uniaxial strain applied along the direction of $\bm{\delta}_{2}$ and $J_{2}^{0}$ is the Heisenberg spin exchange interaction without the strain applied. So $\delta=J_{2}/J_{2}^{0}$ describes the change proportion of $J_{2}$ under uniaxial strain. Phenomenally speaking, we think that when the distance between atoms is reduced by the uniaxial strain, the overlap of electron cloud between atoms increases, which results in an increase of $|J_{2}|$ ($\delta>1$). When the distance between atoms is increased by the uniaxial strain, the overlap of electron cloud between atoms decreases, which results in an decrease of $|J_{2}|$ ($\delta<1$). So we use $\delta$ to represent the uniaxial strain along the direction of $\bm{\delta}_{2}$. And we make the assumption that only the Heisenberg spin exchange interaction along $\bm{\delta}_{2}$ changes, without lattice deformation [30]. And in our work, we take $g_{J}\mu_{B}B=-0.1$ meV, and $\mathcal{K}=-0.2$ meV.
@@
Figure 6: The direction of magnon intrinsic second-order Hall heat current under the temperature gradient along the direction of y-axis. (a) An uniaxial strain in the tensile direction ($\delta<1$) is applied to the system. (b) An uniaxial strain in the compression direction ($\delta>1$) is applied to the system.
@@
Then we study the effects of strain and changing DM strength on $\kappa_{xyy}$. In Fig. 5 (a) and (b), we plot $\kappa_{xyy}$ curve with respect to temperature under different strain and different DM strength respectively. As the temperature rises, $\kappa_{xyy}$ increases rapidly at low temperature, and decreases after reaching a maximum value. There is no $\kappa_{xyy}$ without strain. $\kappa_{xyy}$ with strain of $\delta=0.8$ and $\delta=1.2$ have different sign. And the $\kappa_{xyy}$ curve increases with the increasing of DM strength. As shown in Fig. 5 (c), we can reverse the direction of intrinsic second-order magnon Hall heat current by changing the direction of the strain. And we can increase the intrinsic second-order magnon Hall thermal conductivity by increasing the strength of strain. As shown in Fig. 5 (d), the intrinsic second-order magnon thermal Hall conductivity increases with the increasing of DM strength.
In Fig. 6, we plot a schematic diagram of the direction of intrinsic second-order Hall heat current of magnon when a temperature gradient along the direction of y-axis is applied. When we apply the tensile uniaxial strain ($\delta<1$) along the $\bm{\delta}_{2}$ direction, the second-order intrinsic Hall heat current of magnon flows in the positive direction of x-axis. When we apply the compression uniaxial strain ($\delta>1$) along the $\bm{\delta}_{2}$ direction, the second-order intrinsic Hall heat current of magnon flows in the negative direction of x-axis.
@@
When the second version of the present manuscript is being prepared, we note a similar work from Harsh Varshney et al. in which they calculated the density matrix modification of temperature gradient, then calculated the thermodynamic average value of magnon energy multiplied by velocity and the energy magnetization current through the density matrix [31]. The intrinsic part of heat current has been extracted under the dilute impurity limit (DIL). In our work, we perform extra symmetry analysis in two-dimensional hexagonal ferromagnetic systems. We find that for existence of the non-zero intrinsic nonlinear Hall thermal conductivity of magnon in hexagonal ferromagnet, it is necessary to apply a strain to break the symmetry of $\mathcal{C}_{3}^{z}$. However, in the model calculation by Harsh Varshney et al., intrinsic nonlinear magnon thermal conductivity can be obtained in non-strain hexagonal ferromagnet model, which is different to our results. In addition, we propose the notion of thermal Berry-connection polarizability and work out compact formalism with transparent physical picture, which is out of the scale of Ref. 31. Moreover, we study the effects of strain on the second order intrinsic nonlinear thermal Hall transport. Since the aim of present work is to illustrate the existence of such an intrinsic second-order magnon thermal Hall effect, the magnon-magnon interaction has not been included and is deserved a further study.
@@ Appendix A The detail of the TVP perturbation method
We denote the zero-order perturbed part of $a_{m}(\mathbf{k},t)$ as $a_{m}^{(0)}(\mathbf{k},t)$, and denote the first-order perturbed part of $a_{m}(\mathbf{k},t)$ as $a_{m}^{(1)}(\mathbf{k},t)$. The initial condition is $a_{m}(\mathbf{k},0)=\delta_{mn}$, and the zeroth-order coefficients satisfy $\dot{a}_{m}^{(0)}(\mathbf{k},t)=0$. So we can get $a_{m}^{(0)}(\mathbf{k},t)=\delta_{mn}$. Substituting the zeroth-order coefficients into Eq. (LABEL:A9), we can know that the first-order coefficients satisfy
@@
Then we take the zero frequency limit $\omega\to 0$, and remove non-physical items that increase over time [27]
@@
To get a steady-state, we should introduce a relaxation $\tau$ due to scattering. We take the long-time approximation $(t\gg\tau)$, such that all oscillatory terms drop out [27]. Therefore, we can get the corrected Bloch state
@@ B.0.1 The perturbation from temperature gradient
Analogy the gravitational potential in electron system [32], Hamiltonian under the temperature gradient can be described as
@@ B.0.2 The corrected Bloch state
We follow the method in Ref. 5, and derive the modified Bloch state. Without the perturbation, the wave packet can be expressed as
@@
From [2], we know $\bra{\psi_{l}(\mathbf{k}^{\prime})}\hat{\mathbf{r}}\ket{\psi_{m}(\mathbf{k})}=i\partial_{\mathbf{k}^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})\delta_{lm}+\bm{\mathcal{A}}_{lm}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, so we can get